# Short-Long Convolutions Help Hardware-Efficient Linear Attention to Focus on Long Sequences

### v2

@@ 2.2 Linear Recurrent Models
Continuous form. A model with a continuous time latent space transforms a one-dimensional input signal $u(t)$ into a latent state $x(t)$ that has $d_{s}$ dimensions, and then this $x(t)$ is transformed into a one-dimensional output signal $y(t)$. Concretely, we can define the process:
@@
Eq. 3 is utilized in prior research to model long sequences. For instance, Gu et al. (2020a) suggest a set of matrices called HiPPO (high-order polynomial projection operators) to initialize $\mathbf{A}$. The HiPPO matrices are constructed to enable the state $x(t)$ at time $t$ to remember the history of the input $u(t)$ up to the present.
Discrete time state space model.
@@
Gu et al. (2021) develop the S4 model to efficiently compute Eq. 5. Specifically, $\mathbf{C}$ in Eq. 3 is randomly initialized, and $\mathbf{A}$ and $\mathbf{B}$ are initialized as
@@ 4.3 CHELA Model
Linformer (Wang et al., 2020a) 35.70 53.94 52.27 38.56 76.34 51.36 5.5$\times$
@@ 7 Related Works
There are various attempts to reduce the quadratic time and space complexity of vanilla attention. One approach is to employ sparse attention. That is, each token only attends to a subset of all the tokens according to pre-defined patterns, e.g., neighboring tokens within a fixed-size window. Some examples include Sparse Transformer (Child et al., 2019), BlockBERT (Qiu et al., 2020), Longformer (Beltagy et al., 2020), ETC (Ainslie et al., 2020), BigBird (Zaheer et al., 2020), HEPOS (Huang et al., 2021), and Poolingformer (Zhang et al., 2021). Another approach is to use low-rank projection (Wang et al., 2020a). Similar methods include Nyströmformer  (Xiong et al., 2021), Synthesizer (Tay et al., 2021), Transformer-LS (Zhu et al., 2021), and Luna  (Ma et al., 2021). However, these approaches face difficulty when handling causal tasks, such as auto-regressive language modeling. Kernel-based approaches can be used to approximate the full attention $\mathrm{Attn}(\mathbf{X})$. In these approaches, the quadratic-time softmax attention is replaced by fast linear-time kernel approximations (e.g., Gaussian and arc-cosine kernel). Some examples include Linear Transformer (Katharopoulos et al., 2020), Performer (Choromanski et al., 2020), Random Feature Attention (Peng et al., 2021), and FMMformer (Nguyen et al., 2021). Both low-rank projection and kernel-based approaches approximate the full attention, and thus, they often suffer from non-negligible approximation errors. We can adopt clustering-based approaches, where we divide $\mathbf{Q}$ or $\mathbf{K}$ into several clusters, and perform inter-cluster attention. Such methods include Reformer (Kitaev et al., 2020), Clusterformer (Wang et al., 2020b), Sinkhorn Transformer (Tay et al., 2020a), Fast Transformer (Vyas et al., 2020), Routing Transformer (Roy et al., 2021), and FLASH (Hua et al., 2022).

### v3

@@ 2.2 Linear Recurrent Models
Continuous formulation. A model with a continuous time latent space transforms a one-dimensional input signal $u(t)$ into a latent state $x(t)$ that has $d_{s}$ dimensions, and then this $x(t)$ is transformed into a one-dimensional output signal $y(t)$. Clearly, we can define the process:
@@
Eq. 3 is utilized in prior research to model long sequences. For instance, Gu et al. (2020a) suggests a set of matrices called HiPPO (high-order polynomial projection operators) to initialize $\mathbf{A}$. The HiPPO matrices are constructed to enable the state $x(t)$ at time $t$ to remember the history of the input $u(t)$ up to the present.
Discrete formulation.
@@
Gu et al. (2021) devised the S4 model to effectively calculate Eq. 5. In particular, the initialization of $\mathbf{C}$ in Eq. 3 is random, and both $\mathbf{A}$ and $\mathbf{B}$ are initialized as
@@ 4.3 CHELA Model
Linformer (Wang et al., 2020) 35.70 53.94 52.27 38.56 76.34 51.36 5.5$\times$
@@ 7 Related Works
A variety of efforts have been made to decrease the quadratic time and space complexity of standard attention mechanisms. One method is to utilize “sparse attention,” where each token only attends to a subset of all the tokens based on predefined patterns, such as neighboring tokens within a fixed-size window. (Child et al., 2019) started the attempt to sparse the attention, and then there were a lot more followers, such as ETC (Ainslie et al., 2020), Longformer (Beltagy et al., 2020), BigBird (Zaheer et al., 2020), Poolingformer (Zhang et al., 2021), and HEPOS (Huang et al., 2021) are some examples of this approach. Another option is to utilize “low-rank projection,” as mentioned in the work by (Wang et al., 2020). Similar techniques include Nyströmformer (Xiong et al., 2021), Synthesizer (Tay et al., 2021), and Luna  (Ma et al., 2021). However, these methods encounter challenges when dealing with causal tasks, such as auto-regressive language modeling. Another approach uses “clustering method,” where we partition $\mathbf{Q}$ or $\mathbf{K}$ into multiple clusters and perform inter-cluster attention. Examples of such methods include Sinkhorn Transformer (Tay et al., 2020a), Reformer (Kitaev et al., 2020), Routing Transformer (Roy et al., 2021), and simplified FLASH (Hua et al., 2022), etc. “Methods based on kernels” can be utilized to approximate the complete attention $\mathrm{Attn}(\mathbf{X})$. These methods replace the quadratic-time softmax attention with fast linear-time kernel approximations (such as Gaussian and arc-cosine kernels). Some instances of this approach include Linear Transformer (Katharopoulos et al., 2020), Performer (Choromanski et al., 2020), and FMMformer (Nguyen et al., 2021), etc. Both low-dimensional projection and methods based on kernels are employed to estimate full attention and, as a result, are vulnerable to significant approximation errors.